OpenStudy (anonymous):

will give medal. Find the points on the curve y = tan(x), -π/2 < x < π/2, where the tangent is parallel to the line y = 2x. Give all values correct to 3 decimal places.

OpenStudy (anonymous):

your answer has a bunch of symbols so i cant see the answer..

OpenStudy (anonymous):

im sorry, its doing it again :/

OpenStudy (anonymous):

yeah 1 sec sorry

OpenStudy (anonymous):

y=tan(x) y=sec^2(x) sec^2(x)=2 sec(x)=+- sqrt 2 cos(x)=+- 1sqrt2 so answers is (pi/4,1) (-pi/4_1)

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

how is sec(x)= =-sqrt 2 ?

OpenStudy (anonymous):

or + sqrt 2

OpenStudy (anonymous):

i submitted my answers and it said they were wrong @n648c788

OpenStudy (loser66):

the answer is correct, just typo there, \(\dfrac{\sqrt 2}{2}\) @lalalastudy how can it be wrong?

OpenStudy (anonymous):

im not sure, they were all wrong except for the 1 in the first ordered pair

OpenStudy (loser66):

y' = sec^2 x =\(\dfrac{1}{cos^2x}\) and this y' =2 (because the line // y = 2x)

OpenStudy (anonymous):

i mean the 1 in the second ordered pair

OpenStudy (loser66):

if 1 is correct, then -pi/4 must be correct , also , because if x = -pi/4, then tan x = 1 so as pi/4

OpenStudy (anonymous):

alright i got them right!

OpenStudy (anonymous):

thank you